OpenStudy (anonymous):

How would I determine if this series converges or diverges: sum from 1 to infinity of (sin(1/n^2))

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \sin(\frac{ 1 }{ n^2 })\]

OpenStudy (anonymous):

sin(1/n^2 )

OpenStudy (turingtest):

use\[\sin x<x\]on \((0,\infty)\)

OpenStudy (anonymous):

so sin(1/x^2) is less than 1/x^2 for all x > 0, 1/x^2 is a convergent p-series, thus sin(1/x^2) converges by Comparison Test?

OpenStudy (turingtest):

yep :)

OpenStudy (anonymous):

nice! thanks :)

OpenStudy (turingtest):

welcome!